OpenStudy (anonymous):

(a) A circle is divided into 6 sectors in such a way that the angles of the sectors are in arithmetic progression. The angle of the largest sector is 4 times the angle of the smallest sector. Giventhat the radius of the circle is 5 cm, find the perimeter of the smallest sector. [6] (b) The first, second and third terms of a geometric progression are 2k + 3, k + 6 and k, respectively. Given that all the terms of the geometric progression are positive, calculate (i) the value of the constant k, [3] (ii) the sum to infinity of the progression. [2]

OpenStudy (somy):

okay do you know what an arithmetic progression looks like

OpenStudy (anonymous):

yea

OpenStudy (somy):

okay so if the 6th sector is 4 times bigger than the first sector what must be the progression factor?

OpenStudy (somy):

for the 6 sectors lets say the first sector was a angles a,a+r,a+2r,a+3r...., now a+a+r+a+2r+a+3r...a+5r=360 and u know that a+5r=4a

OpenStudy (popitree):

let the angels be=a, a+b, a+2b,a+3b,a+4b and a+5b now a+5b=4a =>b=3a/5; now total angle in circle = 360 so, 6a+15b = 360 a=24 or a = 2pi/15; now perimeter = 2r+ar/pi =2r+2r/15 =32r/15 =96cm

OpenStudy (somy):

2 equations 2 unknowns solve for a and r, and the he rest